OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 6. Find (f0x)(-5)

OpenStudy (anonymous):

f of x, where x = -5. f(-5) = -2(-5)-7 solve for f(-5)

OpenStudy (jdoe0001):

ahemm did you mean -> f( g(x) ) ?

OpenStudy (anonymous):

^that was what I was wondering

OpenStudy (anonymous):

i think question is fog(x).

OpenStudy (jdoe0001):

or ( f o g)(-5) ?

OpenStudy (anonymous):

( f o g )(-5) ^^

OpenStudy (jdoe0001):

gyanu looks "foggy" yes

OpenStudy (jdoe0001):

\(\bf f(x) = -2x - 7 \qquad {\color{red}{ g(x)}} = -4x + 6 \\ \quad \\ (f\circ g)(x)\implies f(\quad g(x)\quad )=-2({\color{red}{ -4x + 6}}) - 7 \\ \quad \\ (f\circ g)({\color{blue}{ 15}})\implies f(\quad g({\color{blue}{ 15}})\quad )=-2({\color{red}{ -4({\color{blue}{ 15}}) + 6}}) - 7\)

OpenStudy (jdoe0001):

shoot.. is -5.. ahemm ok well \(\bf f(x) = -2x - 7 \qquad {\color{red}{ g(x)}} = -4x + 6 \\ \quad \\ (f\circ g)(x)\implies f(\quad g(x)\quad )=-2({\color{red}{ -4x + 6}}) - 7 \\ \quad \\ (f\circ g)({\color{blue}{ -5}})\implies f(\quad g({\color{blue}{ -5}})\quad )=-2({\color{red}{ -4({\color{blue}{ -5}}) + 6}}) - 7\)

OpenStudy (anonymous):

f of g of x, where x=-5. f(g(-5)): f(gx) = f(x), where g(x) = x f(gx) = -2(-4x+6)-7 f(gx) = 8x-12-7 f(gx) = 8x-19 Let x = -5 8(-5)-19 = -59 or ============ f(g(x)), where x = -5 g(-5) = -4(-5)+6 g(-5) = 26 f(26) = -2(26)-7 f(26) = 59